Please help me to find the correct answer to f(x) = (x - 3)^2 when f(2d), I tried the following f(2d) = (2d - 3)^2 = 4d^2 + 9, I think I am missing something. I also have one other f(x) = (x - 3)^2 when f(c + 6), I tried the following f(c + 6) = (c + 6 - 3)^2 = c^2 + 36 + 9 = c^2 + 45, I do not think it is correct. Please help me solve these two functions. Thank you for your time.

Answer by rothauserc(4718)

f(x) = (x - 3)^2 = x^2 - 6x +9
f(2d) = (2d)^2 -6(2d) +9
f(2d) = 4d^2 -12d -9 = (2d -3)^2